# EKF-Based Radar-Inertial Odometry with Online Temporal Calibration

### v1

@@ IV-D Measurement Update with Radar Measurements
where $N$ is the total number of detected points, $v_{d}^{i}$ denotes the Doppler velocity of the $i$-th point, ${}^{R}\mathbf{p}_{f}^{i}$ is the position vector of the $i$-th point in the radar frame, and ${}^{R}\mathbf{v}_{R}$ denotes the radar ego-velocity. To estimate the radar ego-velocity ${}^{R}\mathbf{v}_{R}$ from noisy radar measurements, various methods, such as RANSAC and m-estimator-based optimization, have been proposed. In this work, we adopt the 3-point RANSAC-LSQ [11], a simple yet robust method that efficiently eliminates outliers and estimates the radar ego-velocity ${}^{R}\mathbf{v}_{R}$, which is used in the measurement update.
In the case where the IMU and radar are rigidly connected, the radar ego-velocity measurement model can be expressed using the system state. As derived in [11], the radar ego-velocity is expressed as:
@@
In the measurement update, the residual $\mathbf{r}$ is computed as the difference between the radar ego-velocity ${}^{R}\mathbf{v}_{R}$, estimated from the radar measurements, and the predicted radar ego-velocity ${}^{R}\hat{\mathbf{v}}_{R}$ from the state. The residual is expressed as:
@@
As illustrated in Fig. 1, $t^{\prime}=t+t_{d}$ represents the IMU measurement time, which serves as the reference time stream in the filter, while $t$ represents the radar measurement time used in the measurement update after being aligned to the IMU time stream. The term $\mathbf{n}_{r}$ denotes the noise of the measurement. The function $\mathbf{h}(\tilde{\mathbf{x}})$ is a nonlinear function that relates the state error $\tilde{\mathbf{x}}$ to the radar ego-velocity measurement residual. For use in the EKF, this function is linearized with respect to the system state. The measurement Jacobian matrix $\mathbf{H}$ is expressed as follows:
@@
$=$ $\mathbf{H}_{q}\left({}^{G}\hat{\mathbf{R}}_{I}\left(t^{\prime}\right)\left({}^{I}\bm{\omega}_{I}^{m}\left(t^{\prime}\right)-\hat{\mathbf{b}}_{g}\left(t^{\prime}\right)\right)\right)$
@@
Each time new radar measurement is received, the measurement update is performed based on the IMU time stream.
@@ IV-E Online Temporal Calibration
The proposed method estimates the time offset between the IMU and the radar in real-time by employing the radar ego-velocity measurement model. By accounting for the time offset, the proposed method ensures that both propagation and measurement updates are performed based on a common time stream, ensuring that the measurements from both sensors are synchronized.
The time offset is propagated using a noise model $n_{d}$, as described in Eq. (7). If the time offset is constant over time or approximately known, it can be estimated without a noise model. However, the time offset varies across sensor models, making it difficult to predefine in most cases. Furthermore, when the vehicle is moving slowly, the impact of the time offset becomes less significant, making it harder to estimate. For this reason, the time offset is modeled as a random walk.
In the measurement model presented in Eq. (13), the factors affected by the temporal misalignment between sensors are not only the state variables ${}^{G}\mathbf{R}_{I}$, ${}^{G}\mathbf{v}_{I}$, and $\mathbf{b}_{g}$, but also the gyroscope measurement ${}^{I}\bm{\omega}_{I}^{m}$. Although $\mathbf{b}_{g}$, which does not change significantly over time, is negligible, failing to account for the time offset causes the state to propagate over a misaligned time stream, leading to errors in the estimates of ${}^{G}\mathbf{R}_{I}$ and ${}^{G}\mathbf{v}_{I}$. Moreover, the improper use of the gyroscope measurement further degrades the estimation accuracy, and this error accumulates over time. Therefore, accounting for the time offset is crucial to maintain the accuracy and consistency of the state estimates.
@@ V-A Experimental Setup
We conduct experiments using three datasets, comprising a total of 15 sequences. One is our self-collected dataset, captured with a handheld platform as shown in Fig. 2, while the other two are public radar datasets: ICINS2021 [13], and ColoRadar [21]. The sensors on our platform include a 4D FMCW radar, specifically the Texas Instruments AWR1843BOOST, and an Xsens MTI-670-DK IMU. No additional hardware triggers are used between the sensors, and the sensor data is recorded using an Intel NUC i7 onboard computer. The experiments are conducted in an indoor area equipped with a motion capture system to obtain precise ground truth. The extrinsic calibration between the IMU and the radar is performed manually. To highlight the significance of temporal calibration in RIO, we design the dataset with two levels of difficulty. Sequences 1 to 3 feature standard motion patterns, while Sequences 4 to 7 introduce more rotational motion to induce larger errors due to the time offset, providing a clearer demonstration of its impact.
@@
In [13], the ICINS2021 dataset is collected using a Texas Instruments IWR6843AOP radar sensor, an Analog Devices ADIS16448 IMU sensor, and a camera. A microcontroller board is used for active hardware triggering to accurately capture the timing of the radar measurements. Data is collected using both handheld and drone platforms. The handheld sequences, “carried_1” and “carried_2”, are referred to as “ICINS 1” and “ICINS 2”, while the drone sequences, “flight_1” and “flight_2”, are referred to as “ICINS 3” and “ICINS 4”, respectively. The ground truth is provided through visual-inertial SLAM, which performs multiple loop closures, offering a pseudo-ground truth. In [21], the ColoRadar dataset is collected using a Texas Instruments AWR1843BOOST radar sensor, a Microstrain 3DM-GX5-25 IMU sensor, and a LiDAR mounted on a handheld platform. No specific synchronization setup is used between the sensors. The sequences, “arpg_lab_run0” and “arpg_lab_run1”, are referred to as “ColoRadar 1” and “ColoRadar 2”, while the sequences “ec_hallways_run0” and “ec_hallways_run1” are referred to as “ColoRadar 3” and “ColoRadar 4”, respectively. The ground truth is generated via LiDAR-inertial SLAM, which includes loop closures, offering a pseudo-ground truth.
@@ V-B Evaluation
For the performance comparison, the open-source EKF-RIO [11], which uses the same measurement model but does not account for temporal calibration, is employed. All parameters are kept identical to ensure a fair comparison. In the proposed method, the time offset $t_{d}$ is initialized to 0.0 seconds for all sequences, reflecting a typical scenario where the initial time offset is unknown. The experimental results are evaluated using the open-source tool EVO [22]. Figure 3 illustrates the estimated trajectories compared to the ground truth for visual comparison, with one representative result from each dataset. Due to the stochastic nature of the RANSAC algorithm used in radar ego-velocity estimation, the averaged results from 100 trials across all datasets are presented. We compare the root mean square error (RMSE) of both absolute pose error (APE) and relative pose error (RPE), with the RPE calculated at 10-meter intervals.
APE evaluates the overall trajectory by calculating the difference between the ground truth and the estimated poses for all frames, making it particularly useful for assessing the global accuracy of the estimated trajectory. However, APE can be sensitive to significant rotational errors that occur early or in specific sections, potentially overshadowing smaller errors later in the trajectory. In contrast, RPE focuses on local accuracy by aligning poses at regular intervals and calculating the error, allowing discrepancies over shorter segments to be highlighted. When the temporal calibration between sensors is not accounted for, errors can accumulate over time, making RPE evaluation essential. Both metrics offer valuable insights, providing a comprehensive evaluation of the trajectory.
V-B1 Self-Collected Dataset
@@
As shown in Table V-B, error values are analyzed with fixed offsets set at 0.05-second intervals for both Sequence 1 and Sequence 4, which feature different motion patterns. The results show that the time offset falls within the -0.10 to -0.15 second range, where the highest accuracy in terms of APE and RPE is observed for both sequences. The proposed method, which utilizes online temporal calibration, estimates the time offset as -0.114 seconds for Sequence 1 and -0.115 seconds for Sequence 4, closely matching the range found through fixed offset testing. In both cases, the proposed method achieves improved performance in terms of both APE and RPE, demonstrates its effectiveness in accurately estimating the time offset.
@@
(Trajectory Length)
@@
Ours (-0.113 s) 0.809 2.388 0.074 0.870
Since the radar delay is generally larger than IMU delay, the time offset $t_{d}$, representing the difference between these delays, typically takes a negative value. To evaluate the robustness of the estimation, different initial values of $t_{d}$ ranging from 0.0 to -0.3 seconds are tested. Figure 4 illustrates the estimated time offset for each initial setting, along with the 3-sigma boundaries. As $t_{d}$ is estimated from radar ego-velocity, it cannot be determined while the platform is stationary. Once the platform starts moving, the filter begins estimating $t_{d}$ and quickly converges to a stable value. The filter converges to a stable time offset of -0.114 ± 0.001 seconds in Sequence 1 and -0.115 ± 0.001 seconds in Sequence 4.
Table V-B1 presents the performance comparison between the proposed method with online temporal calibration and EKF-RIO across seven sequences. The proposed method outperforms EKF-RIO, significantly reducing both APE and RPE across all sequences. Specifically, it reduces APE translation error by an average of 56%, APE rotation error by 75%, RPE translation error by 50%, and RPE rotation error by 58% compared with EKF-RIO. Despite using the same measurement model, the performance improvement is achieved solely by applying propagation and updates based on a common time stream through the proposed online temporal calibration.
Figure 4: Time offset estimation with 3-sigma boundaries for different initial values in Sequence 1 and 4.
@@
LiDAR-Inertial [23] Velodyne VLP-32 0.006 s
Visual-Inertial [18] PointGrey Bumblebee2 0.047 s
Radar-Inertial TI AWR1843BOOST 0.113 s
V-B2 Open Datasets
Table V-B2 presents the results from the two open datasets. The ICINS dataset incorporates a hardware trigger for the radar, which we use to validate the accuracy of the time offset estimation for the proposed method. In this setup, a microcontroller sends radar trigger signals, prompting the radar to begin scanning. The radar data is timestamped based on the actual trigger signal, providing a pseudo-ground truth for time offset estimation. Theoretically, if the sensors are time-synchronized through triggers, the time offset $t_{d}$ is expected to be close to 0.0 seconds. The proposed method estimates the time offset to be an average of 0.016 ± 0.003 seconds. Despite this slight discrepancy, the proposed method demonstrates comparable or improved performance on average in both APE and RPE compared with EKF-RIO. Although the ICINS dataset includes hardware-triggered signals for the radar, there is no such trigger signal for the IMU in the dataset, which may introduce a delay in IMU measurements. As defined in Eq. (1), we attribute the estimated positive time offset to this IMU delay, explaining the difference from the expected value.
@@ V-B2 Open Datasets
The ColoRadar dataset, widely used for performance comparison in the RIO field, is utilized to assess if the proposed method generalizes well across different datasets. As shown in Table V-B2, the proposed method also demonstrates performance improvements over EKF-RIO in terms of both APE and RPE on average. However, the extent of improvement is smaller compared with the self-collected dataset, which can be explained by differences in trajectory characteristics. The radar ego-velocity model utilizes not only the accelerometer but also the gyroscope measurements. As illustrated in Fig. 3, the ColoRadar dataset involves movement over a larger area with less rotation, leading to a smaller impact of the time offset on performance. Nonetheless, the proposed method achieves 33% reduction in RPE translation error, demonstrating its effectiveness even in this less challenging trajectory. On average, the time offset $t_{d}$ is estimated to be -0.111 ± 0.003 seconds, similar to the time offset found in the self-collected dataset. This consistency is likely due to the use of the same radar sensor model in both datasets, further validating the reliability of the proposed method across different environments.
@@
(Trajectory Length)

### v2

@@ IV-D Measurement Update with Radar Measurements
where $N$ is the total number of detected points, $v_{d}^{i}$ denotes the Doppler velocity of the $i$-th point, ${}^{R}\mathbf{p}_{f}^{i}$ is the position vector of the $i$-th point in the radar frame, and ${}^{R}\mathbf{v}_{R}$ denotes the radar ego-velocity. To estimate the radar ego-velocity ${}^{R}\mathbf{v}_{R}$ from noisy measurements, various methods, such as RANSAC and m-estimator-based optimization, have been proposed. In this work, we adopt the 3-point RANSAC-LSQ [11], a simple yet robust method that efficiently eliminates outliers and estimates the radar ego-velocity ${}^{R}\mathbf{v}_{R}$, which is used in the measurement update.
When the IMU and radar are rigidly connected, the radar ego-velocity measurement model can be expressed using the system state. As derived in [11], the radar ego-velocity is expressed as:
@@
To incorporate temporal calibration into the radar update, the proposed method evaluates the predicted radar ego-velocity at the IMU time stream, offset by the estimated time offset $\hat{t}_{d}$. Accordingly, the residual $\mathbf{r}$ is computed as the difference between the radar ego-velocity ${}^{R}\mathbf{v}_{R}(t)$, estimated from the current radar measurements, and the predicted radar ego-velocity ${}^{R}\hat{\mathbf{v}}_{R}(t^{\prime})$. The residual is expressed as:
@@
Here, $t$ denotes the radar measurement timestamp, and $t^{\prime}=t+\hat{t}_{d}$ represents the corresponding IMU time, which serves as the reference time stream in the filter. The term $\mathbf{n}_{r}$ denotes the noise of the measurement. The function $\mathbf{h}(\tilde{\mathbf{x}})$ is a nonlinear function that relates the state error $\tilde{\mathbf{x}}$ to the radar ego-velocity measurement residual. For use in the EKF, this function is linearized with respect to the system state. The measurement Jacobian matrix $\mathbf{H}$ is expressed as follows:
@@
$=$ $\mathbf{H}_{q}\left({}^{I}\bm{\omega}_{I}^{m}\left(t^{\prime}\right)-\hat{\mathbf{b}}_{g}\left(t^{\prime}\right)\right)$
@@
Each time a new radar measurement is received, the measurement update is performed based on the IMU time stream.
@@ IV-E Online Temporal Calibration
The proposed method estimates the time offset between the IMU and the radar in real-time by employing the radar ego-velocity measurement model. By accounting for the time offset, the proposed method ensures that both propagation and measurement updates are performed based on a common time stream, thereby synchronizing the measurements from both sensors.
The time offset is propagated using a noise model $n_{d}$, as described in Eq. (7). If the time offset is constant over time or approximately known, it can be estimated without a noise model. However, the time offset varies across sensor models, making it difficult to predefine in most cases. For this reason, the time offset is modeled as a random walk.
In the measurement model presented in Eq. (13), the factors affected by the temporal misalignment between sensors are not only the state variables ${}^{G}\mathbf{R}_{I}$, ${}^{G}\mathbf{v}_{I}$, and $\mathbf{b}_{g}$, but also the gyroscope measurement ${}^{I}\bm{\omega}_{I}^{m}$. Although $\mathbf{b}_{g}$, which does not change significantly over time, is negligible, failing to account for the time offset causes the state to be propagated using accelerometer and gyroscope measurements taken at incorrectly shifted times, offset by $t_{d}$, as illustrated in Fig. 1. This temporal misalignment leads to errors in the estimates of ${}^{G}\mathbf{R}_{I}$ and ${}^{G}\mathbf{v}_{I}$. Moreover, using the gyroscope measurement at an incorrect time instant introduces additional errors, and these errors accumulate over time. To correct this, the proposed method estimates the time offset $t_{d}$ from the difference in radar ego-velocity induced by temporal misalignment. This difference can arise in various motion scenarios, such as rapid changes in acceleration or angular velocity. In contrast, when the platform is stationary or exhibits only mild motion, the ego-velocity change is small, making $t_{d}$ difficult to be estimated. The Jacobian in Eq. (16) is derived based on the estimated state and IMU measurements evaluated at the shifted time $t^{\prime}$ and is used for estimating the time offset.
@@ V-A Experimental Setup
We validate the proposed method using both simulated and real-world datasets. In the simulation, we use Gazebo with a drone equipped with a virtual IMU and a radar sensor. To simulate realistic behavior, noise is added to the IMU and radar ego-velocity measurements. A time delay of -0.15 seconds is intentionally applied to the radar timestamps to evaluate the performance of the time offset estimation.
For real-world validation, we evaluate the proposed method on three datasets, comprising a total of 15 sequences. One is our self-collected dataset, captured with a handheld platform as shown in Fig. 2, while the other two are public radar datasets: ICINS2021 [13] and ColoRadar [21]. The sensors on our platform include a Texas Instruments (TI) AWR1843BOOST radar and an Xsens MTI670 IMU. No additional hardware triggers are used between the sensors, and the sensor data is recorded using an Intel NUC i7 onboard computer. The experiments are conducted in an indoor area equipped with a motion capture system to obtain precise ground truth. The extrinsic calibration between the IMU and the radar is performed manually. To highlight the significance of temporal calibration in RIO, we design the dataset with two levels of difficulty. Sequences 1-3 involve smoother motion with smaller gyroscope changes over the time offset interval, while Sequences 4-7 exhibit greater variation, leading to a larger radar ego-velocity discrepancy and clearer impact of the time offset.
@@
In [13], the ICINS2021 dataset was collected using a TI IWR6843AOP radar, an ADIS16448 IMU, and a camera. A microcontroller board was used for active hardware triggering to ensure accurate radar timing. Data was captured on both handheld and drone platforms. The handheld sequences “carried_1” and “carried_2” are referred to as “ICINS 1” and “ICINS 2”, respectively. The drone sequences “flight_1” and “flight_2” are referred to as “ICINS 3” and “ICINS 4”, respectively. Ground truth is provided by visual-inertial SLAM with multiple loop closures, serving as pseudo-ground truth. In [21], the ColoRadar dataset was collected using a TI AWR1843BOOST radar, a 3DM-GX5-25 IMU, and a LiDAR on a handheld platform. No hardware synchronization was applied. The sequences “arpg_lab_run0” and “arpg_lab_run1” are referred to as “ColoRadar 1” and “ColoRadar 2”, respectively, while “ec_hallways_run0” and “ec_hallways_run1” are referred to as “ColoRadar 3” and “ColoRadar 4”, respectively. Ground truth is obtained via LiDAR-inertial SLAM with loop closures.
@@ V-B Evaluation
For the performance comparison, the open-source EKF-RIO [11], which uses the same measurement model but does not account for temporal calibration, is employed. All parameters are kept identical to ensure a fair comparison. In the proposed method, the time offset $t_{d}$ is initialized to 0.0 seconds for all sequences, reflecting a typical scenario where the initial time offset is unknown. The experimental results are evaluated using the open-source tool EVO [22]. Figure 3 illustrates the estimated trajectories compared to the ground truth for visual comparison, with one representative result from each dataset. Due to the stochastic nature of the RANSAC algorithm used in radar ego-velocity estimation, the averaged results from 100 trials across all datasets are presented. We compare the root mean square error (RMSE) of both absolute pose error (APE) and relative pose error (RPE), with the RPE calculated at 10-meter intervals, both evaluated after origin alignment.
V-B1 Simulation
Figure 4: Simulation results: (a) Trajectory comparison, and (b) top: difference in radar ego-velocity due to temporal shift by the time offset, and bottom: time offset estimation under different sensor noise levels.
The simulation experiment aims to validate the performance and importance of the proposed temporal calibration under a known time offset. Figure 4(a) shows the trajectory comparison between EKF-RIO without temporal calibration and the proposed method, demonstrating improved accuracy by correcting temporal misalignment.
As discussed in Section IV-E, the time offset is estimated from the discrepancy in radar ego-velocity caused by temporal misalignment. The upper plot in Fig. 4(b) illustrates this discrepancy by comparing the radar ego-velocity at the true measurement time (i.e., $t-0.15$), the delayed timestamp (i.e., $t$), and the time adjusted by the estimated offset (i.e., $t+\hat{t}_{d}$). A larger discrepancy leads to a stronger correction in the estimated time offset. As the proposed method gradually compensates for the offset, this difference diminishes over time, indicating improved temporal alignment. The lower plot in Fig. 4(b) illustrates how increased sensor noise affects the stability of time offset estimation. Results with 30×30$\times$30 × IMU noise and 3×3$\times$3 × radar noise are shown for comparison. The estimation becomes more unstable with higher noise levels, especially with radar noise, since the method depends not only on IMU measurements but also on the estimated state. Accurate radar measurements are crucial, as they directly influence the state estimation in RIO.
V-B2 Self-Collected Dataset
@@
As shown in Table V-B2, error values are analyzed with fixed offsets at 0.05-second intervals for Sequence 1 and Sequence 4, which feature different motion patterns. The best APE and RPE occur within the offset range of -0.10 to -0.15 seconds. The proposed method estimates the offset as -0.114 and -0.115 seconds for Sequence 1 and 4, respectively, closely matching this range. In both cases, it achieves better APE and RPE, validating the accuracy of the online estimation.
@@
(Trajectory Length, $\Delta\omega$)
@@
Ours (-0.113 s) 0.809 2.388 0.074 0.872
*$\Delta\omega$ is defined as the norm of the averaged gyroscope measurement over the interval $[t-0.11,t]$, capturing the rotational motion during that period.
@@
LiDAR-Inertial [23] Velodyne VLP-32 -0.006 s
Visual-Inertial [18] PointGrey Bumblebee2 -0.047 s
Radar-Inertial TI AWR1843BOOST -0.113 s
To evaluate the robustness of the estimation, different initial values of $t_{d}$ ranging from 0.0 to -0.3 seconds are tested. Figure 5 illustrates the estimated time offset for each initial setting, along with the 3-sigma boundaries. As $t_{d}$ is estimated from radar ego-velocity, it cannot be determined while the platform is stationary. Once the platform starts moving, the filter begins estimating $t_{d}$ and quickly converges to a stable value. The filter converges to a stable time offset of -0.114 ± 0.001 seconds in Sequence 1 and -0.115 ± 0.001 seconds in Sequence 4.
Table V-B2 presents the performance comparison between the proposed method with online temporal calibration and EKF-RIO across seven sequences. The proposed method outperforms EKF-RIO, significantly reducing both APE and RPE across all sequences. Specifically, it reduces APE translation error by an average of 56%, APE rotation error by 75%, RPE translation error by 50%, and RPE rotation error by 57% compared with EKF-RIO. Despite using the same measurement model, the performance improvement is achieved solely by applying propagation and updates based on a common time stream through the proposed online temporal calibration.
Figure 5: Time offset estimation with 3-sigma boundaries for different initial values: (a) Sequence 1, and (b) Sequence 4.
Figure 6: Time offset estimation in ICINS 1: (a) estimation results for different artificial offsets, and (b) effect of propagation noise modeling under a delay of $-0.15$s.
V-B3 Open Datasets
The ICINS dataset provides radar data synchronized via a hardware trigger, allowing it to serve as a reference for validating time offset estimation. To evaluate the proposed method, we apply artificial delays to the radar timestamps. As shown in Fig. 6(a), the proposed method estimates the time offsets with an average error, $e_{t_{d}}=|t_{d}^{\text{true}}-\hat{t}_{d}|$, of approximately 0.015 seconds. Figure 6(b) shows the effect of the propagation noise on time offset estimation. Increasing the noise level accelerates convergence but reduces stability, while a smaller noise improves stability at the cost of slower convergence. These results highlight the necessity of noise modeling and emphasize the importance of appropriately tuning the noise parameter to balance responsiveness and robustness in time offset estimation. The results for the ICINS dataset without artificial delay are presented in Table V-B3. The proposed method estimates the time offset as 0.016 ± 0.003 seconds. Despite the slight deviation from the expected value of 0.0 seconds, the method achieves comparable or improved APE and RPE performance compared to EKF-RIO. Although the ICINS dataset includes hardware-triggered signals for the radar, it does not provide such triggering for the IMU. As defined in Eq. (1), the estimated positive time offset is attributed to this IMU delay, explaining the difference from the expected value.
@@ V-B3 Open Datasets
The ColoRadar dataset, widely used for performance comparison in the RIO field, is utilized to assess if the proposed method generalizes well across different datasets. As shown in Table V-B3, the proposed method also demonstrates performance improvements over EKF-RIO in terms of both APE and RPE on average. However, the extent of improvement is smaller compared with the self-collected dataset, which can be explained by differences in motion characteristics. This is attributed to the lower rotational motion present in the ColoRadar dataset, as indicated by smaller values of $\Delta\omega$, resulting in reduced impact of the time offset. Nonetheless, the proposed method achieves 33% reduction in RPE translation error, demonstrating its effectiveness even in this less challenging trajectory. On average, the time offset $t_{d}$ is estimated to be -0.111 ± 0.003 seconds, similar to the time offset found in the self-collected dataset. This consistency is likely due to the use of the same radar sensor model in both datasets, further validating the reliability of the proposed method across different environments.
@@
(Trajectory Length, $\Delta\omega$)
@@
*$\Delta\omega$ is omitted for the ICINS sequences, as they are reported to be time-synchronized.